# ESIQA: Perceptual Quality Assessment of Vision-Pro-based Egocentric Spatial Images

### v1

@@ 2.3.2 Questionnaire Analysis
Post-test questionnaires can help better understand the QoE of egocentric spatial images and design future studies. All questions are shown in Figure 7. For Q1 and Q2, we derive ranking scores for the three viewing modes based on the sorting results, which reflect the overall ranking of the options. The ranking score for a viewing mode is calculated by
$S=\frac{\sum(f\times w)}{n},$ (3)
where $S$ indicates the ranking score for this viewing mode, $f$ represents the frequency of each rank, $w$ denotes the weight assigned to each rank (e.g., the top rank is weighted 3, and the lowest rank is weighted 1), and $n$ is the total number of participants. For Q3 and Q4, we compute the average scores from questionnaire responses to evaluate user dizziness. The statistical results of the replies and some general observations are as follows.
  • Q1: 3D-immersive mode, 3D-window mode, and 2D mode obtain ranking scores of 2.27, 2.18, and 1.55, respectively. The results suggest that most subjects prefer the experience of 3D-immersive and 3D-window modes viewing in Apple Vision Pro compared to the experience of the 2D mode viewing on the screen display, which is consistent with our previous analysis in Section 2.3.1.
  • Q2: We derive the ranking scores for each viewing mode: 3.00, 1.73, and 1.14 for 3D-immersive, 3D-window, and 2D modes, respectively. The results indicate participants generally consider that the 3D-immersive scene offers the most immersive experience.
  • Q3 & Q4: The scores for dizziness in the 3D-immersive and 3D-window viewing modes are 4.27 and 2.27, respectively. This result suggests that while the 3D-immersive mode offers a superior immersive experience, it also induces more significant dizziness compared to the 3D-window mode.
@@ 2.3.3 Database Attributes Analysis
Following the methodologies outlined in [29] and [12], we conduct statistical analysis for our ESIQAD in terms of five low-level vision feature dimensions, including sharpness, brightness, colorfulness, contrast, and cumulative probability of blur detection metric (CPBD). Considering that the ranges of these features are different, we normalized their values to $\left[0,1\right]$ for better illustration. Figure 8 shows the kernel distribution of each feature for our ESIQAD. It can be observed that the spatial images in our database have a wide distribution in brightness, colorfulness, contrast, and CPBD features, showing their extensive diversity. However, our database shows a narrow distribution range for the sharpness feature, notably around 0.3.
@@ 3 Experiment
Table 1: Performance comparison of the state-of-the-art NR IQA models on our ESIQAD under three viewing modes. The best performances are marked in RED, and the second-best performances are marked in BLUE.
Dimension 2D 3D-window 3D-immersive
Model SRCC KRCC PLCC SRCC KRCC PLCC SRCC KRCC PLCC
@@
Resnet18 [13] 0.7574 0.5628 0.7006 0.6120 0.4554 0.6678 0.5127 0.3754 0.5974
Resnet34 [13] 0.6882 0.5107 0.6632 0.6374 0.4839 0.6492 0.6213 0.4582 0.6952
VGG16 [26] 0.6641 0.4747 0.7031 0.5958 0.4436 0.6718 0.5041 0.3666 0.6125
VGG19 [26] 0.6883 0.5022 0.6916 0.6015 0.4470 0.6747 0.5106 0.3731 0.6007
@@
3.1 Benchmark Models
To evaluate the performance of various existing IQA models on predicting human visual perception for egocentric spatial images, we employ 22 state-of-the-art no-reference (NR) IQA models for comparison, which are designed for traditional 2D images. The selected models can be categorized into two groups:
@@ 3.1 Benchmark Models
    Hand-crafted models, which include BRISQUE [20], CORNIA [34], QAC [32], BMPRI [18], NIQE [21], ILNIQE [35], HOSA [31], FISBLIM [11], LPSI [30], BPRI-PSS [17], BPRI-LSSs [17], BPRI-LSSn [17], and BPRI [17].
@@
    Deep learning-based models, including CNNIQA [14], Resnet18 [13], Resnet34 [13], VGG16 [26], VGG19 [26], HyperIQA [27], WaDIQaM [2], MANIQA [33], and TReS [9].
For traditional hand-crafted IQA models, we apply these models to directly predict the quality scores of the left and right views of an egocentric spatial image, respectively. The predicted score from the left view, which is displayed in the 2D viewing mode, directly serves as the 2D quality score for the image. For both 3D-window and 3D-immersive modes, the quality score for the egocentric spatial image is calculated by averaging the scores from the left and right views.
For deep learning-based IQA models, we train and test these models on our ESIQAD. We partition the database into training and testing sets with the ratio of 4:1. The training parameters are set the same as those in the officially released version. Considering that these deep learning-based IQA models are designed for traditional 2D images, we feed the left view of spatial images into the models to predict quality scores for the 2D viewing mode. Since spatial images cannot be directly input into these models, we fine-tune the architectures of these deep learning-based models to accommodate a dual-image (6 channels) input, comprising both left and right views of the spatial image. These fine-tuned models are then utilized to predict quality scores under 3D-window and 3D-immersive viewing modes.
3.2 Performance Analysis
We evaluate the performance of these benchmark models using three evaluation criteria, i.e., Spearman rank correlation coefficient (SRCC), Pearson linear correlation coefficient (PLCC), and Kendall’s rank correlation coefficient (KRCC). These three metrics are utilized to measure the prediction monotonicity. Before computing these criteria, predicted scores are normalized using a five-parameter logistic function [25]:

### v2

@@ 1.3.2 Questionnaire Analysis
Post-test questionnaires can help better understand the QoE of egocentric spatial images and design future studies. All questions are shown in Fig. 7. For Q1 and Q2, we derive ranking scores for the three display modes based on the sorting results, which reflect the overall ranking of the options. The ranking score for a display mode is calculated by
$\vspace{-4pt}S=\frac{\sum(f\times w)}{n},\vspace{-4pt}$ (3)
where $S$ indicates the ranking score for this display mode, $f$ represents the frequency of each rank, $w$ denotes the weight assigned to each rank (e.g., the top rank is weighted 3, and the lowest rank is weighted 1), and $n$ is the total number of participants. For Q3 and Q4, we compute the average scores from questionnaire responses to evaluate user dizziness. The statistical results of the replies and some general observations are as follows: (1) Q1: 3D-immersive mode, 3D-window mode, and 2D mode obtain ranking scores of 2.27, 2.18, and 1.55, respectively. The results suggest that most subjects prefer the experience of 3D-immersive and 3D-window modes viewing in Apple Vision Pro compared to the experience of the 2D mode viewing on the screen display, which is consistent with our previous analysis in Section 1.3.1. (2) Q2: We derive the ranking scores for each display mode: 3.00, 1.73, and 1.14 for 3D-immersive, 3D-window, and 2D modes, respectively. The results indicate participants generally consider that the 3D-immersive scene offers the most immersive experience. (3) Q3 & Q4: The scores for dizziness in the 3D-immersive and 3D-window display modes are 4.27 and 2.27, respectively. This result suggests that while the 3D-immersive mode offers a superior immersive experience, it also induces more significant dizziness compared to the 3D-window mode.
@@ 1.3.3 Database Attributes Analysis
Following the methodologies outlined in [21] and [47], we conduct statistical analysis for our ESIQAD in terms of five low-level vision feature dimensions, including sharpness, brightness, colorfulness, contrast, and cumulative probability of blur detection metric (CPBD). Considering that the ranges of these features are different, we normalized their values to $\left[0,1\right]$ for better illustration. Fig. 8 shows the kernel distribution of each feature for our ESIQAD. It can be observed that the spatial images in our database have a wide distribution in brightness, colorfulness, contrast, and CPBD features, showing their extensive diversity. However, our database shows a narrow distribution range for the sharpness feature due to the fixed resolutions.
2 Proposed Method
2.1 Overall Pipeline
[…]
2.2 Visual State Space Duality Block
The visual state space duality (VSSD) block develops from the state space model (SSM), known for its efficiency in modeling sequential data [18], which can be formulated as:
$h(t)=Ah(t-1)+Bx(t),\quad y(t)=Ch(t),$ (4)
where $A$, $B$, and $C$ are learned parameters defining the system’s state transitions. Mamba2 [6] leverages both the linear recurrence of SSM and the quadratic dual form, introducing state space duality (SSD), which is based on block decompositions of semiseparable matrices, To enhance the SSD block in Mamba2 for vision applications, the visual state space duality (VSSD) block utilizes a bidirectional scanning strategy to eliminate causality, replacing the SSD with an NC-SSD [42]. Compared with SSD, VSSD has implemented many improvements. Specifically, the causal convolution 1D in the NC-SSD block is replaced with a depth-wise convolution (DWConv) with a kernel size of 3. Aligning with classical vision transformers [9, 30], the VSSD block also integrates a feed-forward network (FFN) after the NC-SSD to improve channel communication. Additionally, a local perception unit (LPU) [20] is added before the NC-SSD block and FFN to improve local feature detection.
Figure 10: Visualization of input images and their corresponding heatmaps obtained from different stages of VSSD blocks, which are derived by averaging the vector $\textbf{{V}}^{\textit{l}}_{i}$ across various channels.
In our model, we employ multi-stage VSSD blocks to hierarchically extract spatial features from binocular views of egocentric spatial images. The architecture of the VSSD block is depicted in Fig. 9(b). In the VSSD block, the input X is processed through linear layers and convolutions to obtain the parameters $[A,B,C,\Delta,Z]$, which define the state transition and output behavior of the system. The VSSD process is simply formulated as follows:
$\vspace{-2pt}\operatorname{NC-SSD}(\textbf{{X}})=\operatorname{LinearAttn}(\Delta,A,B,C)\cdot Z,\vspace{-2pt}$ (5)
$^{\prime}=\textbf{{X}}+\operatorname{NC-SSD}(\textbf{{X}}),\vspace{-2pt}$ X ) , (6)
$\vspace{-2pt}\tilde{\textbf{{X}}}=\textbf{{X}}^{\prime}+\operatorname{FFN}(\operatorname{LN}(\textbf{{X}}^{\prime})),\vspace{-2pt}$ (7)
[…]
To illustrate the effectiveness of the multi-stage VSSD blocks intuitively, we visualize the feature maps obtained from different stages of VSSD blocks as heatmaps in Fig. 10 (using the $\textbf{{V}}^{\textit{l}}_{i}$ as an example). The results indicate that the multi-stage VSSD blocks progressively extract semantic features of the input image, allowing the model to focus on elements that are most critical for quality assessment.
2.3 Attention mechanisms
2.3.1 Cross Attention Block
Cross attention mechanisms are widely used in various vision tasks due to their ability to effectively capture relationships between different inputs [5, 11, 27]. In our model, we leverage cross attention to mimic the human brain’s cognitive preference process of integrating visual information from both eyes to perceive egocentric spatial images. This mechanism simulates how the brain forms preferences and creates an immersive experience of these images under 3D modes. Specifically, we use the cross attention module to fuse features extracted from the left and right views, capturing the disparity information between them to derive interactive features.
The cross attention block works by computing interactions between a query from one view and keys from the other view. Specifically, for the left and right view features $\textbf{{V}}^{l}_{i}$ and $\textbf{{V}}^{r}_{i}$, the cross attention module generates query (Q), key (K), and value (V) projections through linear transformations of the input features. The query from one view interacts with the keys of the other view via a dot-product operation to produce an attention map that reflects the correlations between the views, where the attention map A can be formulated as:
$\textbf{A}=\operatorname{Softmax}\left(\frac{\textbf{{Q}}^{l}\left(\textbf{{K}}^{r}\right)^{T}}{\sqrt{d_{k}}}\right),$ (8)
where $\textbf{{Q}}^{l}$ and $\textbf{{K}}^{r}$ are the query and key projections from the left and right views, respectively, and $d_{k}$ is the scaling factor. This attention map is then used to weight the value projection $\textbf{{V}}^{r}_{i}$, aggregating information from the right view into the left view’s context.
2.3.2 Transposed Attention Block
Traditional self attention employs key-query dot-product interactions to establish global connections among spatial patches but overlooks channel-wise information [46]. To address this issue, the transposed attention block applies self attention across channels rather than the spatial dimension [51, 53]. This method calculates cross-covariance across channels to produce an attention map that implicitly encodes global context.
In our model, we utilize the transposed attention block to prioritize the channels of the integrated spatial features $\textbf{{F}}_{i}$ and enhance downstream local interactions by leveraging the globally contextual information encoded by the attention map. Specifically, the transposed attention generates query (Q), key (K), and value (V) projections via three independent linear projections of $\textbf{{F}}_{i}$. Then we reshape the query and key projections for their dot-product interaction, producing a transposed attention map $\textbf{A}\in\mathbb{R}^{c^{\prime}\times c^{\prime}}$. The architecture of the transposed attention block is depicted in Fig. 9(c), which can be formulated as:
$\textbf{A}=\operatorname{Softmax}\left(\frac{\textbf{{Q}}\textbf{{K}}}{\sqrt{d_{k}}}\right),$ (9)
where $d_{k}$ is the dimension of each head in the attention mechanism. Then we obtain the transposed attention output feature $\tilde{\textbf{{F}}}_{i}$ by applying the residual connection with $\textbf{{F}}_{i}$:
$\tilde{\textbf{{F}}}_{i}=\textbf{{W}}_{p}\cdot(\textbf{{V}}\cdot\textbf{A})+\textbf{{F}}_{i},$ (10)
where $\textbf{{W}}_{p}$ represents a trainable linear projection matrix.
2.4 Quality Regression
The quality regression module uses a multi-layer perceptron (MLP) to compute the final quality scores, as shown in Fig. 9(a). This module consists of 7 layers, i.e., three linear layers for dimensionality reduction, two ReLU activations for non-linearity, and two dropout layers to prevent overfitting. Before inputting to the MLP, we apply average pooling to extract the channel-wise average feature of $\tilde{\textbf{{F}}}_{i}$, generating a channel-level global representation. We then concatenate the global representations across different semantic levels. The quality regression process is formulated as:
$\hat{Q}=\operatorname{MLP}\left(\operatorname{avg}(\tilde{\textbf{{F}}}_{1})\oplus\operatorname{avg}(\tilde{\textbf{{F}}}_{2})\oplus\cdots\oplus\operatorname{avg}(\tilde{\textbf{{F}}}_{n})\right),$ (11)
where $\oplus$ is the concatenation operator and the $\operatorname{avg}$ denotes the average pooling operation.
@@ 3 Experiment
Table 1: Detailed model specifications of the ESIQAnet variants, categorized as Micro (M), Tiny (T), Small (S), and Base (B). The specifications include the number of blocks at each stage, the channels used, attention heads, parameter counts (#Param.), and floating point operations per second (FLOPs).
Model Blocks Channels Heads #Param. FLOPs
ESIQAnet-M [2, 2, 8, 4] [48, 96, 192, 384] [2, 4, 8, 16] 77.57M 7.84G
ESIQAnet-T [2, 4, 12, 4] [64, 128, 256, 512] [2, 4, 8, 16] 92.78M 14.74G
ESIQAnet-S [3, 4, 21, 5] [64, 128, 256, 512] [2, 4, 8, 16] 104.71M 18.90G
ESIQAnet-B [3, 4, 21, 5] [96, 192, 384, 768] [3, 6, 12, 24] 155.01M 38.64G
Table 2: Performance comparison of the state-of-the-art NR IQA models and the four ESIQAnet variants on our ESIQAD under three display modes. In the top-5 results for each criteria, the best result is marked in RED, the second-best result is marked in BLUE, and the remaining three are marked in Underlined.
Mode 2D 3D-window 3D-immersive
Method SRCC KRCC PLCC SRCC KRCC PLCC SRCC KRCC PLCC
@@
CNNIQA-v16 [43] 0.6641 0.4747 0.7031 0.5958 0.4436 0.6718 0.5041 0.3666 0.6125
CNNIQA-v19 [43] 0.6883 0.5022 0.6916 0.6015 0.4470 0.6747 0.5106 0.3731 0.6007
CNNIQA-r18 [22] 0.7574 0.5628 0.7006 0.6120 0.4554 0.6678 0.5127 0.3754 0.5974
CNNIQA-r34 [22] 0.6882 0.5107 0.6632 0.6374 0.4839 0.6492 0.6213 0.4582 0.6952
@@
ESIQAnet-M 0.8334 0.6533 0.8683 0.8297 0.6476 0.8638 0.7633 0.5771 0.7980
ESIQAnet-T 0.8427 0.6622 0.8814 0.8212 0.6384 0.8557 0.7739 0.5876 0.8129
ESIQAnet-S 0.8518 0.6766 0.8895 0.8314 0.6488 0.8610 0.7797 0.5927 0.8138
ESIQAnet-B 0.8515 0.6721 0.8900 0.8375 0.6531 0.8639 0.7912 0.6038 0.8180
3.1 Experimental Setup
3.1.1 Compared Methods
To evaluate the performance of our proposed model, we employ 22 state-of-the-art no-reference (NR) IQA models for comparison, which are designed for traditional 2D images. The selected models can be categorized into two groups:
@@ 3.1.1 Compared Methods
    Hand-crafted models: BRISQUE [34], CORNIA [52], QAC [50], BMPRI [33], NIQE [35], ILNIQE [54], HOSA [49], FISBLIM [19], LPSI [48], BPRI-PSS [32], BPRI-LSSs [32], BPRI-LSSn [32], and BPRI [32].
@@
    Deep learning-based models: CNNIQA [23], CNNIQA-r18 [22], CNNIQA-r34 [22], CNNIQA-v16 [43], CNNIQA-v19 [43], HyperIQA [44], WaDIQaM [4], MANIQA [51], and TReS [16].
3.1.2 Evaluation Metrics
We evaluate the performance of these IQA models using three evaluation criteria, i.e., Spearman rank correlation coefficient (SRCC), Kendall’s rank correlation coefficient (KRCC), and Pearson linear correlation coefficient (PLCC). These three metrics are utilized to measure the prediction monotonicity. Before computing these criteria, predicted scores are normalized using a five-parameter logistic function [41]:
@@ 3.1.2 Evaluation Metrics
Additionally, we utilize receiver operating characteristic (ROC) analysis [24, 25] as an additional evaluation method for IQA metrics. This method assesses two key aspects, i.e., whether two samples distinctly differ in quality and, if so, which one is superior. Through “Different vs. Similar ROC Analysis", we determine whether various objective metrics can distinguish image pairs with and without significant qualitative differences. We classify image pairs with significant differences into pairs with positive and negative differences. The “Better vs. Worse ROC Analysis" is applied to test if various objective metrics can distinguish images with positive and negative differences. The area under the ROC curve (AUC) values of the two analyses is reported in this paper, where higher values represent better performance.
Table 3: Performance comparison of the state-of-the-art NR IQA models and the proposed ESIQAnet-B on the two subsets of our ESIQAD under three display modes. Bold entries in RED, BLUE, and BLACK are the best, second-best, and third-best performances, respectively.
Subset of Captured Images
Mode 2D 3D-window 3D-immersive
Method SRCC KRCC PLCC SRCC KRCC PLCC SRCC KRCC PLCC
BMPRI [33] 0.1707 0.1192 0.0737 0.1358 0.0968 0.2759 0.1224 0.0866 0.3030
BPRI [32] 0.1795 0.1269 0.1147 0.1914 0.1332 0.1660 0.1513 0.1052 0.1235
BPRI-LSSn [32] 0.3699 0.2566 0.6240 0.2822 0.1944 0.5286 0.2723 0.1856 0.5093
BPRI-LSSs [32] 0.3512 0.2429 0.4265 0.3001 0.2115 0.4003 0.2705 0.1897 0.3745
BPRI-PSS [32] 0.1385 0.0927 0.3379 0.1003 0.0673 0.2709 0.1260 0.0851 0.2656
BRISQUE [34] 0.3644 0.2517 0.6014 0.2624 0.1809 0.5025 0.2493 0.1672 0.4766
CORNIA [52] 0.2040 0.1370 0.4603 0.2173 0.1484 0.3917 0.1692 0.1127 0.3633
FISBLIM [19] 0.2372 0.1606 0.6093 0.1988 0.1345 0.5148 0.1912 0.1277 0.5048
HOSA [49] 0.2869 0.1952 0.4741 0.2154 0.1448 0.4312 0.2211 0.1482 0.4325
NIQE [35] 0.3719 0.2568 0.6039 0.2917 0.1996 0.5162 0.2788 0.1885 0.5043
ILNIQE [54] 0.3408 0.2351 0.4758 0.2801 0.1909 0.4235 0.2550 0.1734 0.4010
LPSI [48] 0.3588 0.2488 0.6274 0.2752 0.1880 0.5303 0.2655 0.1804 0.5110
QAC [50] 0.1351 0.0886 0.3667 0.0917 0.0590 0.2907 0.0935 0.0604 0.2749
CNNIQA [23] 0.5425 0.3734 0.5296 0.4285 0.2956 0.4379 0.4006 0.2842 0.3969
CNNIQA-v16 [43] 0.5125 0.3857 0.5498 0.4447 0.3165 0.5884 0.4341 0.3095 0.5428
CNNIQA-v19 [43] 0.5580 0.3905 0.5732 0.4634 0.3297 0.6159 0.4494 0.3146 0.5687
CNNIQA-r18 [22] 0.5611 0.3962 0.6006 0.4418 0.3139 0.5571 0.4115 0.2943 0.5166
CNNIQA-r34 [22] 0.5522 0.4089 0.5627 0.4530 0.3399 0.4225 0.4237 0.3196 0.3852
WaDIQaM [4] 0.5067 0.3553 0.5342 0.4595 0.3212 0.5839 0.3992 0.2751 0.5237
HyperIQA [44] 0.6681 0.4968 0.6341 0.5714 0.4082 0.5634 0.4462 0.3101 0.4625
MANIQA [51] 0.7107 0.5234 0.8010 0.7099 0.5190 0.7041 0.6556 0.4880 0.7432
TReS [16] 0.6091 0.4225 0.6618 0.5608 0.3829 0.5990 0.5424 0.3793 0.5881
ESIQAnet-B 0.7690 0.5895 0.8107 0.7406 0.5609 0.7969 0.6907 0.5085 0.7300
Subset of Synthesized Images
Mode 2D 3D-window 3D-immersive
Method SRCC KRCC PLCC SRCC KRCC PLCC SRCC KRCC PLCC
BMPRI [33] 0.4273 0.2962 0.5067 0.3175 0.2178 0.4564 0.3412 0.2259 0.4706
BPRI [32] 0.1052 0.0715 0.3629 0.0562 0.0372 0.2133 0.0124 0.0081 0.2556
BPRI-LSSn [32] 0.4280 0.2929 0.4762 0.4331 0.3018 0.4232 0.4477 0.3115 0.4336
BPRI-LSSs [32] 0.1990 0.1313 0.2795 0.1041 0.0671 0.2249 0.1282 0.0857 0.2363
BPRI-PSS [32] 0.1011 0.0723 0.3269 0.0223 0.0182 0.1320 0.0175 0.0093 0.1774
BRISQUE [34] 0.2608 0.1677 0.3113 0.3713 0.2545 0.4232 0.3820 0.2465 0.4416
CORNIA [52] 0.2130 0.1333 0.4492 0.1168 0.0804 0.4264 0.1497 0.1055 0.4318
FISBLIM [19] 0.0872 0.0541 0.4282 0.0389 0.0222 0.3712 0.0618 0.0408 0.3248
HOSA [49] 0.4051 0.2816 0.4648 0.4112 0.2877 0.4241 0.4024 0.2731 0.4374
NIQE [35] 0.4328 0.3164 0.5342 0.4569 0.3204 0.4822 0.4855 0.3430 0.5056
ILNIQE [54] 0.1460 0.1030 0.2174 0.1935 0.1329 0.2214 0.1608 0.1127 0.2066
LPSI [48] 0.3952 0.1180 0.3327 0.3720 0.2400 0.3212 0.3833 0.2400 0.3263
QAC [50] 0.2410 0.1604 0.3027 0.1140 0.0735 0.2116 0.0822 0.0590 0.2108
CNNIQA [23] 0.3286 0.2474 0.3671 0.2789 0.2368 0.2734 0.2256 0.1789 0.2269
CNNIQA-v16 [43] 0.4256 0.3133 0.5685 0.4004 0.2797 0.5336 0.3936 0.2747 0.4538
CNNIQA-v19 [43] 0.4413 0.2987 0.5732 0.4594 0.2941 0.5126 0.4278 0.2789 0.3970
CNNIQA-r18 [22] 0.4429 0.3842 0.4806 0.3098 0.2737 0.3902 0.3541 0.2579 0.3669
CNNIQA-r34 [22] 0.4854 0.3358 0.5030 0.3946 0.2494 0.4016 0.3675 0.2438 0.3922
WaDIQaM [4] 0.4977 0.3615 0.5385 0.4746 0.3601 0.4716 0.4487 0.3025 0.4275
HyperIQA [44] 0.4624 0.3632 0.4934 0.4331 0.3158 0.4786 0.4346 0.2947 0.3660
MANIQA [51] 0.7081 0.5323 0.7380 0.6976 0.5006 0.7819 0.6537 0.4670 0.6687
TReS [16] 0.6070 0.4263 0.5795 0.5131 0.3681 0.6354 0.4738 0.3293 0.5833
ESIQAnet-B 0.7534 0.5895 0.7861 0.7370 0.5575 0.7983 0.6863 0.5032 0.6545
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 11: ROC analysis results of 11 outstanding IQA benchmark methods and the ESIQAnet on the ESIQAD in the 3D-immersive (first row) and the 3D-window (second row) display modes. Note that a white/black square in the significance figures means the row metric is statistically better/worse than the column one. A gray square means the row method and the column method are statistically indistinguishable.
3.1.3 Implementation Details
For traditional hand-crafted benchmarks IQA models, we apply these models to directly predict the quality scores of the left and right views of an egocentric spatial image, respectively. The predicted score from the left view, which is displayed in the 2D mode, directly serves as the 2D quality score for the image. For both 3D-window and 3D-immersive modes, the quality score for the egocentric spatial image is calculated by averaging the scores from the left and right views. For deep learning-based benchmarks IQA models, we train and test these models on our ESIQAD. We partition the database into training and testing sets with a ratio of 4:1. The training parameters are set the same as those in the officially released version. Considering that these deep learning-based IQA models are designed for traditional 2D images, we feed the left view of spatial images into the models to predict quality scores for the 2D display mode. Since spatial images cannot be directly input into these models, we fine-tune the architectures of these deep learning-based models to accommodate a dual-image (6 channels) input, comprising both left and right views of the spatial image. These fine-tuned models are then utilized to predict quality scores of spatial images under 3D-window and 3D-immersive display modes.
For our ESIQAnet, we extract multi-scale features after three VSSD stages with $N_{1}$, $N_{2}$, and $N_{3}$ VSSD blocks respectively, and after the final stage with $N_{4}$ MSA blocks. We design four architecture variants for ESIQAnet with different parameters of VSSD and MSA blocks, which are shown in Table 1. When training models for different display modes, we use their corresponding MOSs as labels. Notably, for training in the 2D display mode, we use only the left view of the spatial image as input and omit the cross attention module, while other components of the model remain unchanged. We use the same training and test sets as the benchmark models and exclusively use the training set to train the models and obtain the optimal models that yield the lowest loss. These optimal models are then evaluated on the testing set.
3.2 Performance Analysis
Table 4: Ablation study on the attention mechanisms our ESIQAnet-B. TAB refers to the transposed attention block, CAB refers to our cross attention block, and MSAB indicates the multi-head self attention block. Best performances are indicated with bold.
Mode 2D 3D-window 3D-immersive
# CAB TAB MSAB SRCC KRCC PLCC SRCC KRCC PLCC SRCC KRCC PLCC
1 0.8040 0.6244 0.8312 0.7898 0.6063 0.8252 0.7544 0.5665 0.7988
2 $\checkmark$ 0.8221 0.6361 0.8663 0.8167 0.6100 0.8388 0.7642 0.5791 0.8039
3 $\checkmark$ 0.8230 0.6330 0.8510 0.8023 0.6182 0.8285 0.7556 0.5698 0.8023
4 $\checkmark$ $\checkmark$ 0.8515 0.6721 0.8900 0.8230 0.6384 0.8490 0.7782 0.5904 0.8128
5 $\checkmark$ - - - 0.8083 0.6250 0.8451 0.7742 0.5850 0.8045
6 $\checkmark$ $\checkmark$ - - - 0.8281 0.6403 0.8559 0.7828 0.5918 0.8137
7 $\checkmark$ $\checkmark$ - - - 0.8236 0.6428 0.8502 0.7824 0.5946 0.8063
8 $\checkmark$ $\checkmark$ $\checkmark$ - - - 0.8375 0.6531 0.8639 0.7912 0.6038 0.8180
Table 5: Ablation study on the backbone of our ESIQAnet-B. The best performances are indicated with bold. The second-best performances are marked in Underlined.
Mode 2D 3D-window 3D-immersive
Backbone SRCC PLCC SRCC PLCC SRCC PLCC
ResNet-18 [22] 0.8162 0.8369 0.7938 0.8322 0.7315 0.7808
ResNet-34 [22] 0.8141 0.8471 0.7907 0.8466 0.7251 0.7703
ResNet-50 [22] 0.8224 0.8525 0.7986 0.8326 0.7408 0.7969
ViT [8] 0.8309 0.8601 0.8198 0.8527 0.7552 0.8031
SwinT [30] 0.8290 0.8522 0.8040 0.8340 0.7431 0.7913
VSSD 0.8515 0.8900 0.8375 0.8639 0.7912 0.8180
3.2.1 Evaluation on the ESIQAD
We evaluate the aforementioned 22 state-of-the-art models and our proposed model on the ESIQAD. The experimental results are demonstrated in Table 2. The results show that the four variants of ESIQAnet outperform all benchmarks across all three display modes, demonstrating excellent prediction ability and strong generalization capabilities in different display modes. Specifically, the ESIQAnet-B achieves a 2.78% improvement in SRCC and a 1.47% improvement in PLCC over the best benchmark model MANIQA in the 2D display mode, a 4.82% improvement in SRCC and a 2.73% increase in PLCC in the 3D-window mode, and a 5.33% improvement in SRCC and a 3.70% rise in PLCC in the 3D-immersive mode. From Table 2, we observe that the performance of ESIQAnet improves with the increased number of parameters and computational complexity. More blocks enable the extraction of deeper visual semantic features, while more attention heads in the multi-head self attention mechanism help focus on various image regions simultaneously, enhancing information fusion for a more comprehensive quality assessment. For the subsequent experiments, we use the ESIQAnet-B for evaluation.